# Enhance Eye Disease Detection using Learnable Probabilistic Discrete Latents in Machine Learning Architectures

### v2

@@ Abstract
Ocular diseases, including diabetic retinopathy and glaucoma, present a significant public health challenge due to their high prevalence and potential for causing vision impairment. Early and accurate diagnosis is crucial for effective treatment and management. In recent years, deep learning models have emerged as powerful tools for analysing medical images, such as retina imaging. However, challenges persist in model relibability and uncertainty estimation, which are critical for clinical decision-making. This study leverages the probabilistic framework of Generative Flow Networks (GFlowNets) to learn the posterior distribution over latent discrete dropout masks for the classification and analysis of ocular diseases using fundus images. We develop a robust and generalizable method that utilizes GFlowOut integrated with ResNet18 and ViT models as the backbone in identifying various ocular conditions. This study employs a unique set of dropout masks - none, random, bottomup, and topdown - to enhance model performance in analyzing these fundus images. Our results demonstrate that our learnable probablistic latents significantly improves accuracy, outperforming the traditional dropout approach. We utilize a gradient map calculation method, Grad-CAM, to assess model explainability, observing that the model accurately focuses on critical image regions for predictions. The integration of GFlowOut in neural networks presents a promising advancement in the automated diagnosis of ocular diseases, with implications for improving clinical workflows and patient outcomes.
@@
1 Introduction
The world faces considerable challenges in terms of eye care. Research indicates that in 2020, the estimated global cases of age-related macular degeneration stood at 196 million, and this figure is anticipated to escalate to 288 million by 2040 Wong et al. (2014). According to the World Health Organization, in its report of "World Report on Vision", more than 2.2 billion people suffer from vision impairment or blindness. Importantly, it is estimated that over 1 billion of these cases could potentially have been avoided with proper prevention or effective treatment WTO (2019). The World Vision Report indicates that primary causes of blindness include Glaucoma, Age-Related Macular Degeneration, and Diabetic Retinopathy. Diagnosing these conditions typically involves an ophthalmologist evaluating a patient’s symptoms, analyzing various eye and retina images, and conducting a manual examination. This process is thorough but can be time-consumingWTO (2019). Other researchers highlighted that the prevalence of Age-Related Macular Degenerations (AMDs) is notably higher in Africa and the Eastern Mediterranean regions compared to other areas of the world Xu et al. (2020). The lack of and unequal distribution of medical resources means that preventable and treatable cases of blindness and low vision predominantly affect people in less developed countries and regions. Vision impairment stems from various factors, notably the retina, which is a key element in disorders like glaucoma, diabetic retinopathy, and age-related macular degeneration. Properly addressing eye health requires not only accurate diagnosis but also effective prevention and treatment strategies for these conditions Yang et al. (2021).
Ophthalmology heavily depends on imaging for diagnosis, as the majority of eye conditions are identified through image analysis. However, traditional screening involves handling large volumes of data, is highly subjective, and requires complex data analysis. This presents a significant burden for both patients and ophthalmologists, complicating long-term follow-ups Besenczi et al. (2016). The incorporation of artificial intelligence(AI), particularly machine learning and deep learning, into this field has significantly boosted the efficiency of clinical eye specialists. AI technology processes and analyzes ophthalmic images, thereby streamlining diagnostic procedures Padhy et al. (2019); Yu et al. (2018); Rajpurkar et al. (2022); Zhou et al. (2023). Currently, there has been considerable research on artificial intelligence-assisted diagnosis in diseases such as glaucoma, diabetic retinopathy, retinopathy of prematurity, and age-related macular degeneration (AMD) Ting et al. (2018). However, we found the majority of the models primarily focuse on diagnosing a single ophthalmic disease Li et al. (2021). There are multiple works shown that deep learning algorithms are promising in the diagnosing diabetic retinopathy through retinal fundus image grading Oh et al. (2021); Wang et al. (2022); Son et al. (2020). However, the high performance of these methods often comes with a significant increase in time complexity. Additionally, their performance can be limited by using uniform image sizes, leading to less robust classifications Li et al. (2022).
Along with these issues, a significant limitation of current deep neural networks is their tendency to exhibit overconfidence in predictions and lack a mechanism for capturing uncertainty, particularly when there is a shift in the data distribution between training and testing datasets Folgoc et al. (2021). This issue is especially prominent in medical imaging, where variability in data can impact diagnostic accuracy. While methods such as standard dropout exist to address this, they often fail to capture the multi-modality of posterior distributions over dropout masks. To mitigate these challenges, GFlowOut Liu et al. (2023) has been recently proposed, leveraging Generative Flow Networks (GFlowNets) Bengio et al. (2023) to model the posterior distribution over dropout masks. However, its potential in real-world medical applications remains underexplored. In this work, we bridge this gap by applying GFlowOut to neural networks for ocular disease classification using the Ocular Disease Intelligent Recognition (ODIR) dataset. Our key contribution is demonstrating the utility of GFlowOut in improving model uncertainty estimation and diagnostic accuracy across a diverse set of ocular conditions, providing a robust solution to variability in medical imaging datasets.
@@ 2.1 Generative Flow Networks
Generative Flow Networks (GFlowNets) have recently emerged as a compelling framework for generating complex, high-dimensional objects by modeling the flow of probability through sequences of states. GFlowNets address the challenge of sampling objects in proportion to a predefined reward function by adopting a control problem formulation, where objects are constructed sequentially via probabilistic steps. This methodology enables GFlowNets to efficiently explore and sample from multimodal distributions, making them particularly well-suited for applications requiring diverse and high-quality solutions, such as drug discovery and protein design Bengio et al. (2023).
The versatility of GFlowNets has been demonstrated across various domains, including drug discovery Bengio et al. (2021a), biological sequence design Jain et al. (2022), robust combinatorial optimization Zhang et al. (2022), causal discovery Deleu et al. (2022), and neural network structure learning Pan et al. (2023a). Foundational work has highlighted the ability of GFlowNets to generalize effectively to complex distributions and reduce gradient variance relative to traditional policy gradient methods, thereby establishing a robust framework for probabilistic modeling Malkin et al. (2023b); Bengio et al. (2021b).
Subsequent advancements have further extended the capabilities of GFlowNets. For instance, Pan et al. (2023b) introduced Stochastic GFlowNets to address the challenges posed by stochastic environments, incorporating intrinsic exploration rewards to enhance training. Additionally, Deleu et al. (2022); Nishikawa-Toomey et al. (2022) applied GFlowNets to the generative modeling of discrete and composite objects, with a particular focus on Bayesian structure learning of complex causal graphs. The framework has also been leveraged in approximate maximum-likelihood training of energy-based models, as demonstrated by Zhang et al. (2022), without the need for a predefined target reward. Moreover, GFlowNets have been applied to tackle NP-hard combinatorial optimization problems, providing a promising approach to these computationally intensive tasks Zhang et al. (2022). In the realm of biological sequence design, Jain et al. (2022) employed GFlowNets within an active learning loop to optimize sequence generation. Furthermore, Zimmermann et al. (2023) offered a variational perspective on GFlowNets by formulating variational objectives through the use of KL divergences. Collectively, these studies underscore the adaptability and potential of GFlowNets in addressing a wide array of generative modeling challenges across diverse fields.
@@ 2.2 GFlowOut - Dropout with Generative Flow Networks
Liu et al. (2023) introduced GFlowOut, a novel solution to the challenges inherent in traditional dropout techniques used within neural networks. These challenges include the multi-modality of the posterior distribution over dropout masks and the difficulty in fully utilizing sample-dependent information and the correlation among dropout masks. GFlowOut leverages the principles of Generative Flow Networks (GFlowNets) to enhance dropout regularization by learning the posterior distribution over dropout masks. Traditional dropout methods often struggle to accurately capture the posterior due to the multimodal and discrete nature of dropout masks Liu et al. (2023); Jain et al. (2022).
GFlowOut addresses these limitations by employing GFlowNets to generate layer-wise dropout masks that are conditioned on previous layer activations and labels, thus improving the estimation of uncertainty and robustness to distributional shifts. Empirical evaluations have demonstrated that GFlowOut significantly outperforms standard methods, such as Random Dropout and Contextual Dropout, across a variety of tasks, including image classification under deformations, visual question answering, and real-world clinical predictions Liu et al. (2023). By utilizing the Trajectory Balance objective during training, GFlowOut ensures that the generated masks are proportionate to the reward function, providing a robust framework for improving posterior estimation and effectively leveraging sample-dependent information in neural networks. This results in enhanced generalization and superior performance in downstream tasks Liu et al. (2023); Malkin et al. (2023a).
@@ 3.3 Eye Disease Dataset
Figure 2: Sample images from the ODIR dataset. The top row displays left and right eye fundus photographs of a normal individual, i.e., a person not diagnosed with any ocular disease. The bottom row shows left and right eye fundus photographs of a patient with diabetes.
Table 1: Summary of diseases in the Ocular Disease Intelligent Recognition (ODIR) dataset
Disease Class Count
Normal 2873
Diabetes 1608
Glaucoma 284
Cataract 293
Age-related Macular Degeneration 266
Hypertension 128
Pathological Myopia 232
Other diseases/abnormalities 708
The Ocular Disease Intelligent Recognition (ODIR) dataset Maranhão (2020) is a comprehensive ophthalmic database consisting of records from 5,000 patients, including age information, color fundus photographs of both eyes, and diagnostic keywords provided by medical professionals. This dataset reflects a "real-life" patient cohort, collected by Shanggong Medical Technology Co., Ltd., from multiple hospitals and medical centers across China. The fundus images in the database were captured using various commercially available cameras, such as Canon, Zeiss, and Kowa, leading to variations in image resolution.
The dataset categorizes images into several classes: normal, diabetes, glaucoma, cataract, age-related macular degeneration, hypertension, pathological myopia, and other diseases. For the purposes of this study, we focused on images labeled as normal and diabetes. These images were randomly shuffled and then split into training and testing datasets with an 80% and 20% allocation, respectively. Due to data constraints, we limited our study to these two classes, though the methodology employed can be extended to other diseases and multi-class classification problems. The pixel values of the images were normalized to lie within the range $[0,1]$.
@@
Prior to being fed into the model, the images underwent several pre-processing steps. Initially, the images were cropped to a size of $224\times 224\times 3$ pixels. They were then resized to $256\times 256\times 3$ pixels, normalized using means of $\mu=[0.485,0.456,0.406]$ and standard deviations of $\sigma=[0.229,0.224,0.225]$, and finally, bi-linear interpolation was applied. These pre-processing steps are consistent with the standard procedures for preparing inputs to ResNet and Vision Transformer models.
For experiments involving noise, Gaussian, Salt, and Speckle noise were added following the completion of the image pre-processing steps. These types of noise were introduced to simulate various noise conditions that could occur in clinical settings, where noisy data might be provided as input to the model. This approach aims to mimic such real-world conditions.

### v3

@@ Abstract
Ocular diseases, including diabetic retinopathy and glaucoma, present a significant public health challenge due to their high prevalence and potential for causing vision impairment. Early and accurate diagnosis is crucial for effective treatment and management. In recent years, deep learning models have emerged as powerful tools for analysing medical images, such as retina imaging. However, challenges persist in model relibability and uncertainty estimation, which are critical for clinical decision-making. This study leverages the probabilistic framework of Generative Flow Networks (GFlowNets) to learn the posterior distribution over latent discrete dropout masks for the classification and analysis of ocular diseases using fundus images. We develop a robust and generalizable method that utilizes GFlowOut integrated with ResNet18 and ViT models as the backbone in identifying various ocular conditions. This study employs a unique set of dropout masks - none, random, bottomup, and topdown - to enhance model performance in analyzing these fundus images. Our results demonstrate that our learnable probablistic latents significantly improves accuracy, outperforming the traditional dropout approach. We utilize a gradient map calculation method, Grad-CAM, to assess model explainability, observing that the model accurately focuses on critical image regions for predictions. The integration of GFlowOut in neural networks presents a promising advancement in the automated diagnosis of ocular diseases, with implications for improving clinical workflows and patient outcomes.The source code for all these experiments can be found at https://github.com/anirudhprabhakaran3/gflowout_on_eye_images.
1 Introduction
The world faces considerable challenges in terms of eye care. Ocular diseases represent serious public health challenges due to their widespread prevalence (Zhou et al., 2023) and potential to cause significant vision impairment (Yang et al., 2021). Studies project that in 2040, there will be an estimated 288 million global cases of age-related macular degeneration (Wong et al., 2014). According to the World Health Organization, in its report of titled “World Report on Vision”, more than 2.2 billion people suffer from vision impairment or blindness. Importantly, it is estimated that over 1 billion of these cases could potentially have been avoided with proper prevention or effective treatment. It indicates that primary causes of blindness include Glaucoma, Age-Related Macular Degeneration, and Diabetic Retinopathy. Diagnosing these conditions typically involves an ophthalmologist evaluating a patient’s symptoms, analyzing various eye and retina images, and conducting a manual examination. This process is thorough but can be time-consuming (WTO, 2019). Other researchers have highlighted that the prevalence of Age-Related Macular Degenerations (AMDs) is notably higher in Africa and the Eastern Mediterranean regions compared to other areas of the world (Xu et al., 2020). The lack of and unequal distribution of medical resources means that preventable and treatable cases of blindness and low vision predominantly affect people in less developed countries and regions. Vision impairment stems from various factors, notably the retina, which is a key element in disorders like glaucoma, diabetic retinopathy, and age-related macular degeneration. Properly addressing eye health requires not only accurate diagnosis but also effective prevention and treatment strategies for these conditions (Yang et al., 2021).
Ophthalmology heavily depends on imaging for diagnosis, as the majority of eye conditions are identified through image analysis. However, traditional screening involves handling large volumes of data, is highly subjective, and requires complex data analysis. This presents a significant burden for both patients and ophthalmologists, complicating long-term follow-ups (Besenczi et al., 2016). Early diagnosis and effective management are crucial in preventing these diseases from progressing to more severe stages. Traditional diagnostic methods, which typically involve manual examination of retinal images, are often time-consuming and subject to variability among practitioners. The incorporation of artificial intelligence (AI), particularly machine learning and deep learning, into this field has significantly boosted the efficiency of clinical eye specialists. AI technology processes and analyzes ophthalmic images, thereby streamlining diagnostic procedures (Padhy et al., 2019; Yu et al., 2018; Rajpurkar et al., 2022; Zhou et al., 2023). Since 2016, Google has applied Deep Learning to analyze retinal images for the detection of diabetic retinopathy (Gulshan et al., 2016). Currently, there has been considerable research on artificial intelligence-assisted diagnosis in diseases such as glaucoma, diabetic retinopathy, retinopathy of prematurity, and age-related macular degeneration (AMD) (Ting et al., 2018). The utilized of Deep Learning has shown immense promise in automating the analysis of medical images, providing more consistent and scalable solutions for disease diagnosis. However, we found that most of the models focus primarily on diagnosing a single ophthalmic disease (Li et al., 2021). There are multiple works showing that deep learning algorithms are promising in the diagnosing diabetic retinopathy through retinal fundus image grading (Oh et al., 2021; Wang et al., 2022; Son et al., 2020). However, the high performance of these methods often comes with a significant increase in time complexity. Additionally, issues around the reliability of these models and their capacity to estimate uncertainty continue to present challenges in clinical decision-making (Li et al., 2022).
Along with these issues, a significant limitation of current deep neural networks is their tendency to exhibit overconfidence in predictions and lack a mechanism for capturing uncertainty, particularly when there is a shift in the data distribution between training and testing datasets (Folgoc et al., 2021). This issue is especially prominent in medical imaging, where variability in data can impact diagnostic accuracy. While methods such as standard dropout exist to address this, they often fail to capture the multi-modality of posterior distributions over dropout masks. To mitigate these challenges, GFlowOut (Liu et al., 2023) has been recently proposed, leveraging Generative Flow Networks (GFlowNets) (Bengio et al., 2023) to model the posterior distribution over dropout masks. However, its potential in real-world medical applications remains underexplored. Our key research question is: How can probabilistic dropout masks be leveraged to enhance the performance of deep learning models in ocular disease detection? We explore Generative Flow Networks as a probabilistic framework to address these limitations. Specifically, we model the posterior distribution over discrete dropout masks to improve classification and analysis of ocular diseases using retinal fundus images. We explore Generative Flow Networks as a probabilistic framework to address these limitations. Specifically, we model the posterior distribution over discrete dropout masks to improve classification and analysis of ocular diseases using retinal fundus images.
Figure 1 illustrates the integration of GFlowOut within the Vision Transformer architecture for eye disease detection. The input fundus image is divided into patches, which are processed through the transformer layers to extract feature representations. In place of traditional dropout layers, GFlowOut layers are applied to model the posterior distribution over discrete dropout masks, enhancing the model’s uncertainty estimation and robustness. The output from the GFlowOut layers is passed through multi-layer perceptron (MLP) heads, which map the processed features to the predicted disease class. This architecture demonstrates the use of learnable probabilistic latents to improve model reliability and diagnostic performance, as detailed in Section 3.1.
@@ 1 Introduction
Our key contribution is demonstrating the utility of GFlowOut in improving model uncertainty estimation and diagnostic accuracy across a diverse set of ocular conditions, providing a robust solution to variability in medical imaging datasets. The integration of GFlowOut into neural networks not only improves model reliability and accuracy but also contributes to the development of more reliable diagnostic tools that can assist in clinical workflows, potentially enhancing patient outcomes.
@@ 2.1 Generative Flow Networks
Generative Flow Networks (GFlowNets) have recently emerged as a compelling framework for generating complex, high-dimensional objects by modeling the flow of probability through sequences of states. GFlowNets address the challenge of sampling objects in proportion to a predefined reward function by adopting a control problem formulation, where objects are constructed sequentially via probabilistic steps. This methodology enables GFlowNets to efficiently explore and sample from multimodal distributions, making them particularly well-suited for applications requiring diverse and high-quality solutions, such as drug discovery and protein design (Bengio et al., 2023).
The versatility of GFlowNets has been demonstrated across various domains, including drug discovery (Bengio et al., 2021a), biological sequence design (Jain et al., 2022), robust combinatorial optimization (Zhang et al., 2022), causal discovery (Deleu et al., 2022), and neural network structure learning (Pan et al., 2023a). Foundational work has highlighted the ability of GFlowNets to generalize effectively to complex distributions and reduce gradient variance relative to traditional policy gradient methods, thereby establishing a robust framework for probabilistic modeling (Malkin et al., 2023b; Bengio et al., 2021b).
Subsequent advancements have further extended the capabilities of GFlowNets. For instance, (Pan et al., 2023b) introduced Stochastic GFlowNets to address the challenges posed by stochastic environments, incorporating intrinsic exploration rewards to enhance training. Additionally, (Deleu et al., 2022; Nishikawa-Toomey et al., 2022) applied GFlowNets to the generative modeling of discrete and composite objects, with a particular focus on Bayesian structure learning of complex causal graphs. The framework has also been leveraged in approximate maximum-likelihood training of energy-based models, as demonstrated by (Zhang et al., 2022), without the need for a predefined target reward. Moreover, GFlowNets have been applied to tackle NP-hard combinatorial optimization problems, providing a promising approach to these computationally intensive tasks (Zhang et al., 2022). In the realm of biological sequence design, (Jain et al., 2022) employed GFlowNets within an active learning loop to optimize sequence generation. Furthermore, (Zimmermann et al., 2023) offered a variational perspective on GFlowNets by formulating variational objectives through the use of KL divergences. Collectively, these studies underscore the adaptability and potential of GFlowNets in addressing a wide array of generative modeling challenges across diverse fields.
@@ 2.2 GFlowOut - Dropout with Generative Flow Networks
(Liu et al., 2023) introduced GFlowOut, a novel solution to the challenges inherent in traditional dropout techniques used within neural networks. These challenges include the multi-modality of the posterior distribution over dropout masks and the difficulty in fully utilizing sample-dependent information and the correlation among dropout masks. GFlowOut leverages the principles of Generative Flow Networks (GFlowNets) to enhance dropout regularization by learning the posterior distribution over dropout masks. Traditional dropout methods often struggle to accurately capture the posterior due to the multimodal and discrete nature of dropout masks (Liu et al., 2023; Jain et al., 2022).
GFlowOut addresses these limitations by employing GFlowNets to generate layer-wise dropout masks that are conditioned on previous layer activations and labels, thus improving the estimation of uncertainty and robustness to distributional shifts. Empirical evaluations have demonstrated that GFlowOut significantly outperforms standard methods, such as Random Dropout and Contextual Dropout, across a variety of tasks, including image classification under deformations, visual question answering, and real-world clinical predictions (Liu et al., 2023). By utilizing the Trajectory Balance objective during training, GFlowOut ensures that the generated masks are proportionate to the reward function, providing a robust framework for improving posterior estimation and effectively leveraging sample-dependent information in neural networks. This results in enhanced generalization and superior performance in downstream tasks (Liu et al., 2023; Malkin et al., 2023a).
@@ 3.1 Model Structure
To further clarify the step-by-step implementation of this approach, Algorithm 1 outlines the detailed pipeline for eye disease detection. The process begins with standard image preprocessing, including normalization and cropping to remove irrelevant regions, followed by partitioning images into patches for efficient processing. These patches are then input into the Vision Transformer’s feature extractor, leveraging self-attention to learn rich feature representations. The GFlowOut layers are applied after each Attention-MLP block to model the posterior distribution of discrete dropout masks, improving robustness against overfitting and data distribution shifts. Finally, these regularized features are passed through the classification head to predict disease probabilities for each input image. This algorithm operationalizes the proposed framework by translating the described architecture into actionable steps, showcasing the practical integration of probabilistic regularization techniques like GFlowOut into state-of-the-art vision models.
The GFlowOut layers are implemented as multiple MLPs. $q(z_{i}|x_{i},y_{i};\phi)$ is implemented as a set of multiple MLPs, once for each layer in the model needing dropout. At layer $l$, the dropout probabilities of all units in layer $l$ are estimated, and repeated for each layer in the model. For more implementation details, you can check the original GFlowOut paper (Liu et al., 2023).
1
Input: Retinal images dataset Images
Output: Disease probability P for each input image
Data: Normalized retinal images, patches of images, feature vectors, attention weights
2
3
4Function process_images(Images):
       // Step 1: Preprocessing the retinal images
5       foreach Image in Images do
6             Image $\leftarrow$ Normalize(Image) // Standardize pixel intensity across images
7             Image $\leftarrow$ Crop(Image) // Remove irrelevant portions of the image
8
      // Step 2: Split the images into patches
9       foreach Image in Images do
10             Patches $\leftarrow$ SplitIntoPatches(Image) // Segment image into smaller patches
11
      // Step 3: Pass patches through Vision Transformer feature extractor
12       foreach Patch in Patches do
13             Features $\leftarrow$ ViTFeatureExtractor(Patch) // Extract feature vectors using the Vision Transformer
14
      // Step 4: Pass features through attention heads
15       foreach Feature in Features do
16             AttentionOutput $\leftarrow$ ApplyAttentionHeads(Feature) // Use attention to highlight relevant data
17
      // Step 5: Apply GFlowOut for enhanced regularization
18       foreach AttentionOutput in AttentionOutputs do
19             RegularizedFeature $\leftarrow$ ApplyGFlowOut(AttentionOutput) // Regularize features using learnable dropout
20
      // Step 6: Compute reward and trajectory balance
21       foreach RegularizedFeature in RegularizedFeatures do
[…]
23             TrajectoryBalance $\leftarrow\prod_{t=0}^{T-1}P(a_{t}\mid s_{t})=\frac{R(s_{T})}{Z}$ // Ensure trajectory balance based on reward
24
      // Step 7: Pass data through the classification head
25       foreach RegularizedFeature in RegularizedFeatures do
26             P $\leftarrow$ ClassificationHead(RegularizedFeature) // Compute probability of disease
27
      // Step 8: Output probabilities
28       foreach Image in Images do
29             Print(P) // Display the disease probability for the input image
30
Algorithm 1 Algorithm for Eye Disease Detection Using Vision Transformer and GFlowOut
@@ 3.2 GFlowOut Masks
Generative Flow Networks (GFlowNets) provide a probabilistic framework to sample objects (e.g., dropout masks) in proportion to a reward function. GFlowOut integrates GFlowNets into neural networks to dynamically learn dropout masks, improving model robustness and performance in tasks such as image classification. GFlowNets generate objects by modeling the flow of probabilities through sequential states. Trajectory Balance Objective aligns mask generation with task-specific rewards.GFlowOut’s dynamic mask generation adapts to data variability, improving regularization and robustness, particularly for tasks like image classification.
Flow Consistency Equation :
The flow consistency equation ensures the flow of probabilities into and out of a state $s$ is balanced:
$F(s)=\sum_{a\in\text{Actions}(s)}F(s,a),$
where:
  • $F(s)$ is the total flow into state $s$.
  • $F(s,a)$ is the flow through state $s$ via action $a$.
Trajectory Balance Objective :
The trajectory balance objective ensures that the probability of generating a trajectory $\tau$ is proportional to the reward of its terminal state. For a trajectory $\tau=(s_{0},a_{0},s_{1},\dots,s_{T})$:
$\prod_{t=0}^{T-1}P(a_{t}\mid s_{t})=\frac{R(s_{T})}{Z},$
where:
  • $P(a_{t}\mid s_{t})$ is the policy probability of taking action $a_{t}$ at state $s_{t}$.
  • $R(s_{T})$ is the reward of the terminal state $s_{T}$.
  • $Z$ is the normalization constant (partition function).
Dropout Mask Probability $P(\text{Mask})$ :
GFlowOut learns the posterior distribution over dropout masks. The probability of a specific mask $M$ is proportional to its reward:
$P(M)\propto R(M).$
Posterior Distribution :
The posterior distribution is computed using a softmax function:
$P(\text{dropout}\mid X,H)=\text{softmax}(W\cdot[X,H]),$
where:
  • $X$ is the input data.
  • $H$ is the context from preceding layers.
  • $W$ are learnable weights.
@@ 3.3 Eye Disease Dataset
In this study, we use three datasets - the Ocular Disease Intelligent Recognition (ODIR) dataset (Maranhão, 2020), the Retinal Fundus Multi-Disease (RFMiD) Image dataset (Panchal et al., 2023) and the Joint Shantou International Eye Center (JSIEC) (Cen et al., 2021) dataset. The ODIR and RFMiD datasets are used for training of the model and evaluation, and the JSIEC dataset is used for out-of-distribution (OOD) and uncertainty estimation experiments.
The Ocular Disease Intelligent Recognition (ODIR) dataset is a comprehensive ophthalmic database consisting of records from 5,000 patients, including age information, color fundus photographs of both eyes, and diagnostic keywords provided by medical professionals. The uniqueness of this dataset is in providing paired data for both left and right eyes along with descriptions. However, for our experiments and model architecture, we consider the two fundus images separate, and extract the diagnosis per image from the description. ODIR-5K provides 8 categories of labels: Normal (N), Diabetes (D), Glaucoma (G), Cataract (C), Age-related Macular Degeneration (AMD), Hypertension (H), Myopia (M), and Other diseases/abnormalities (O).
The Retinal Fundus Multi-Disease (RFMiD) Image Dataset is a multi-label classification dataset for fundus images. We use the 2.0 version of the dataset for our study. The dataset contains a total of 3200 cases, with annotations for 45 different ocular diseases. Given the large number of annotation categories in the dataset, some have very few samples, making this dataset also relevant for research into multi-label long-tail problems and issues with limited samples. The Joint Shantou International Eye Center (JSIEC) dataset consists of 1000 fundus images over 39 categories. We use this subset of the bigger dataset, since we are only using the data for OOD evaluation. Table 1 presents this metadata about all the datasets used in brief.
Table 1: Metadata about the datasets used.
Dataset Name Data Files Classes
ODIR 10000 8
RFMiD 3200 45
JSIEC 1000 39
@@
Prior to being fed into the model, the images underwent several pre-processing steps. Initially, the images were center cropped to a size of $224\times 224\times 3$ pixels. These images were then normalized using means of $\mu=[0.485,0.456,0.406]$ and standard deviations of $\sigma=[0.229,0.224,0.225]$. These pre-processing steps are consistent with the standard procedures for preparing inputs to ResNet and Vision Transformer models.
An important aspect that we wish to study is how well our model would perform with actual clinical data. To emulate the noises in medical imaging equipment, we add Gaussian, Salt-and-Pepper and Speckle noise. Gaussian noise can arise due to electronic noise in the imaging sensors or equipment and presents random variations in pixel intensity as fine grain-like interference across the image. Salt-and-Pepper noise causes random white and black pixels, occurring due to transmission errors, sensor defects or corruption due to image storage or transfer. Speckle noise also occurs as grainy, salt-and-pepper-like interference but has a more localised and granular pattern. This usually occurs in random interference or variations during the imaging process. In experiments using noisy data, we manually add these noises to the images after the standard pre-processing steps are completed.
3.4 Out of Distribution Evaluation and Entropy Calculations
Entropy measures the uncertainty or unpredictability of a probability distribution. For a single sample, the entropy $H$ is given by:
$H=-\sum_{i}{p_{i}log(p_{i})}$
where $p_{i}$ is the predicted probability for class $i$.
Higher entropy indicates higher uncertainty, while lower entropy indicates greater confidence in the prediction. We calculate entropy on both the training dataset and the evaluation datasets. This is done to measure how confident the model is in its predictions for data. Calculating entropy on the OOD datasets allows us to evaluate robustness. A well-trained model should show higher entropy for OOD samples, including uncertainty because these samples deviate from the training distribution. It might be overconfident if a model assigns low entropy (high confidence) to OOD samples.
To get a more quantitative idea about the performance of the model, we also compute the Expected Calibration Error (ECE). Model calibration aims to align the predictions of a model with the true probabilities and thereby making sure that the predictions of a model are reliable and accurate. Although ECE is not specifically designed for OOD detection, better calibration on in-distribution (ID) data indirectly enhances OOD detection by ensuring that confidence scores or uncertainty measures, such as entropy, are meaningful. We compute ECE on the ID dataset to assess the quality of the model’s uncertainty estimation and demonstrate that our proposed method, GFlowOut, achieves significantly lower ECE compared to baseline models.
We define the ECE as
$ECE=\sum_{m=1}^{M}\frac{|B_{m}|}{n}|acc(B_{m})-conf(B_{m})|$
where $M$ is the number of bins, $m$ is the bin number, $|B_{m}|$ is the size of the bin, $conf(B_{m})$ is the average estimated probabilities in bin $m$, defined as
$conf(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\hat{p}_{i}$
and $acc(B_{m})$ is the accuracy per bin $m$, defined as
$acc(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}(\hat{y}_{i}=y_{i})$
A lower ECE score indicates that the model is more calibrated towards the actual probabilities.